OpenStudy (anonymous):

Find the approximate area between the x-axis and the graph of the equation y=x squared from x=0 to 3, using the trapezoid method. Use trapezoids of height 1.

OpenStudy (dumbcow):

 Area of trapezoid: A = 1/2*h*(b1+b2) h=1 b1 and b2 are f(x1) and f(x2) so first trapezoid is from x=0 to1, f(0) = 0, f(1) = 1 A = 1/2(0+1) = 1/2 Next one is x=1 to 2 A = 1/2( f(1) + f(2)) = ? and so on ...